OpenStudy (anonymous):

Prove that \(x\) is both rational and an integer.\[\lim_{n\to\infty}\left(\ln n-\frac{n}{\pi\left(n\right)}\right)=x\]The function \(\pi(n)\) counts the number of primes less than or equal to \(n\). I've seen this equation somewhere before, but I can't remember where. IIRC, \(x\) evaluates to \(1\). Mysterious!

OpenStudy (anonymous):

Furthermore, using Mathematica to algorithmically evaluate \(x\) by using arbitrarily large values of \(n\) does not converge to \(1\)!

OpenStudy (anonymous):

* i love to know what is the answer

OpenStudy (anonymous):

i found something similar \[\lim_{n\to\infty}\frac{\pi(n)}{\frac{n}{\ln n}}=1\] http://en.wikipedia.org/wiki/Prime_number_theorem http://mathworld.wolfram.com/PrimeNumberTheorem.html

OpenStudy (anonymous):

It's only necessary that someone links to a proof. I know I've seen this before, somewhere.

OpenStudy (anonymous):

This is called the prime number theorem.

OpenStudy (anonymous):

Oh, wait, I found it in mathworld. Okay, thanks!

OpenStudy (anonymous):

That's right. This is Legendre's constant. That's where I saw it before! http://en.wikipedia.org/wiki/Legendre%27s_constant

OpenStudy (anonymous):

Thanks for hte link, @eliassaab

OpenStudy (anonymous):

yw